OpenStudy (anonymous):

For the given equation, list the intercepts and test for symmetry: y=x^2-x-20

OpenStudy (anonymous):

for x intercept put y =0 and get x for y intercept put x =0 and get y

OpenStudy (anonymous):

i know that part so then I get 0=x^2-x-20..im lost here

OpenStudy (anonymous):

im not sure how to do the x..its been a while.. haha

OpenStudy (anonymous):

0=x^2-x-20 0 = x^2-5x+4x -20 0 = x(x-5) +4(x-5) 0 = (x+4)(x-5) means that x = -4 and x = 5 which are required x- intercepts

OpenStudy (anonymous):

wow.. im lost.. how did you get the -5x and 4x?

OpenStudy (anonymous):

i feel so dumb haha

OpenStudy (irishboy123):

the crude working and the curve itself attached.

OpenStudy (anonymous):

i got it by guess actualy when we factorize an equation by grouping method we do guesses to find these proper numbers

OpenStudy (anonymous):

have you ever done factorization y grouping method???

OpenStudy (anonymous):

*by

OpenStudy (anonymous):

......i dont know.. i havent taken a math class in so long..